### v1

@@ Abstract
The success of the text-guided diffusion model has inspired the development and release of numerous powerful diffusion models within the open-source community. These models are typically fine-tuned on various expert datasets, showcasing diverse denoising capabilities. Leveraging multiple high-quality models to produce stronger generation ability is valuable, but has not been extensively studied. Existing methods primarily adopt parameter merging strategies to produce a new static model. However, they overlook the fact that the divergent denoising capabilities of the models may dynamically change across different states, such as when experiencing different prompts, initial noises, denoising steps, and spatial locations. In this paper, we propose a novel ensembling method, Adaptive Feature Aggregation (AFA), which dynamically adjusts the contributions of multiple models at the feature level according to various states (i.e., prompts, initial noises, denoising steps, and spatial locations), thereby keeping the advantages of multiple diffusion models, while suppressing their disadvantages. Specifically, we design a lightweight Spatial-Aware Block-Wise (SABW) feature aggregator that adaptive aggregates the block-wise intermediate features from multiple U-Net denoisers into a unified one. The core idea lies in dynamically producing an individual attention map for each model’s features by comprehensively considering various states. It is worth noting that only SABW is trainable with about 50 million parameters, while other models are frozen. Both the quantitative and qualitative experiments demonstrate the effectiveness of our proposed Adaptive Feature Aggregation method. The code is available at https://github.com/tenvence/afa/.
Keywords:
Image Generation Diffusion Models Model Ensembling
@@ 1 Introduction
Diffusion models [50, 17] have progressively become the mainstream models for text-guided image generation [35, 39, 48, 46, 5, 62, 10], which treats generation as an iterative denoising task. Recently, the open-source stable diffusion (SD) [46] model has prompted the development and release of numerous powerful diffusion models within the open-source community (e.g., CivitAI [^1] https://civitai.com/). These models are typically fine-tuned on various expert datasets, showcasing diverse denoising capabilities. Leveraging multiple high-quality models to dig out better generations is an important research direction, which has not been extensively studied.
@@
Existing methods leverage multiple diffusion models through the weighted merging of model parameters, which can be called the static method. The weights are usually manually set [56, 33] or automatically searched through enumeration [4]. In contrast, model ensembling, which can be called the dynamic method, often uses dynamic strategies to fuse multiple models at the feature level. Unlike ensembling for classification models [12] that usually work after the output, ensembling for diffusion models typically needs to work for each block. However, the denoising capabilities of models vary not only at different blocks but also at different spatial locations. As illustrated in Fig. 1, we conduct denoising experiments on different prompts with various initial noises. Then, we plot the proportion of wins (i.e., the model with the least error between the predicted noise and the initial noise), for each model in a certain spatial region, to distinguish the denoising capabilities of the models at different states. In other words, different prompts, initial noises, denoising steps, and spatial locations can all have a significant impact on the denoising capabilities of diffusion models. This implies that an adaptive method is needed to ensure that each diffusion model can dominate the generation at its strongest states.
@@
In this paper, we propose a novel Adaptive Feature Aggregation (AFA) method, which dynamically adjusts the contributions of multiple models at the feature level by taking into account various states, such as prompts, initial noises, denoising steps, and spatial locations. Specifically, we design a lightweight Spatial-Aware Block-Wise (SABW) feature aggregator that adaptively aggregates the block-level intermediate features from multiple U-Net denoisers into a unified one. The core idea of adaptive aggregation lies in dynamically producing an individual attention map for each model’s features by comprehensively taking into account the various states such as text prompts, denoising steps, and spatial locations. A noteworthy aspect of AFA is that only SABW is trainable with about 50 million parameters, while all other models are frozen.
@@ 2 Related Work
Early works for text-guided image synthesis leverage Generative Adversarial Networks (GAN) [14] conditioned on text [53, 61, 65, 71]. Based on the success of Transformers [54], many subsequent works reframe text-guided image synthesis as a sequence-to-sequence task [66, 40, 7, 13, 64]. Recently, diffusion models [50, 17] gradually become mainstream, which treats image generation as an iterative denoising task. By injecting text as a condition into the denoising process, many diffusion-based models achieve significant success in text-guided image synthesis [35, 39, 48, 46, 5, 62, 10]. Among them, the Latent Diffusion Model (LDM) [46] performs the diffusion and reverse process in the latent space, instead of the pixel space, which largely reduces the computational burden. Among all the implementations of the Latent Diffusion Model, the stable diffusion model (SD) is the most famous one. Along with its public availability, the open-source community achieves tremendous success, with the emergence of many excellent fine-tuned models. These high-quality models are typically fine-tuned on various expert datasets, showcasing diverse denoising capabilities. In this paper, we aim to leverage multiple such models to achieve stronger text-to-image generation.
2.0.2 Model Merging.
@@
For vision or language understanding tasks, the effectiveness of parameter merging [11, 58, 32, 20, 25, 8, 21] can be interpreted from the perspectives of loss landscape [57, 3, 52] and task arithmetic [19, 36, 67]. However, merging-based methods have not been widely studied for generation tasks, especially diffusion-based generation. Recently, many intuitive merging-based methods have emerged. One of the popular methods is Weighted Merging [56], which manually determines weights to merge each U-Net [47] parameter of multiple SD models. Although the simplicity, Weighted Merging coarsely allocates weight to all U-Net blocks. As an improvement, Merge Block Weighted (MBW) [33] allows for manually setting different merging weights for the parameters of distinct U-Net blocks, which provides a more fine-grained merging strategy. Furthermore, to reduce the reliance on manual weights, autoMBW [4] attempts to automate the merging process, which enumeratively selects the optimal combination of MBW weights by an aesthetic classifier. However, autoMBW is constrained by the performance bottleneck of the aesthetic classifier, and the enumerative selection leads to a huge time consumption to find the optimal settings. Such merging methods can be called the static methods. In this paper, we aim to leverage multiple diffusion models by the dynamic methods, i.e., ensembling.
2.0.3 Model Ensembling.
@@
Model ensembling is an effective method to achieve better performance [70], which has been widely applied in various vision understanding tasks, such as classification [69, 45, 63], regression [34, 43], clustering [55]. While fewer works focus on the ensembling of generative models, because of the complexity of image space. Vision-Aided GAN [24] focuses on GANs [14], which guides the optimization of a target generator by ensembling pretrained models as a loss. MagicFusion [68] focuses on the diffusion models, which fuses the predicted noises of two expert U-Net denoisers to implement specific applications, such as style transferring and object binding. In this paper, we aim to efficiently ensemble multiple diffusion models to achieve generation improvements.
@@ 3.1 Preliminary
As a type of generative model, the diffusion model [50, 17] consists of two processes, which are the diffusion process and the reverse process, respectively. In the diffusion process (i.e., the forward process), the Gaussian noises are iteratively added to degrade the images over $T$ steps until the images become completely random noise. In the reverse process, a trained denoiser is used to iteratively generate images from the sampled Gaussian noise.
@@
$\mathcal{L}_{\text{denoise}}=\mathbb{E}_{\mathbf{x}_{0},\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I}),\mathbf{c},t}\left\|\boldsymbol{\epsilon}_{\theta}\left(\mathbf{x}_{t},\mathbf{c},t\right)-\boldsymbol{\epsilon}\right\|^{2}\ .$ (1)
Among them, $\mathbf{x}_{t}=\sqrt{\alpha_{t}}\mathbf{x}_{0}+\sqrt{1-\alpha_{t}}\boldsymbol{\epsilon}$ is the noisy image at timestep $t\in[1,T]$, where $\alpha_{t}$ is a predefined scalar from the noise scheduler. $\boldsymbol{\epsilon}$ is the added noise. $\boldsymbol{\epsilon}_{\theta}$ is the denoiser with the learnable parameters $\theta$, which predicts the noise to be removed from the noisy image.
LDM stands out as one of the most popular diffusion models. It performs the diffusion and reverse process in the latent space which is encoded by a Variational Auto-Encoder (VAE) [22, 44]. In LDM, the U-Net structure [47] is used for denoiser, which contains three parts of blocks, which are the down-sampling blocks, the middle block, and the up-sampling blocks, respectively. Each down-sampling block is skip-connected with a corresponding up-sampling block. These blocks are composed of Transformer [54] layers and ResNet [15] layers. In the Transformer layers, the cross-attention mechanism is employed to incorporate conditional textual prompts, which plays a crucial role in text-guided image generation.
Figure 2: Framework of ensembling multiple diffusion models by our Adaptive Feature Aggregation (AFA) method.
@@ 3.2 Adaptive Feature Aggregation
Our proposed Adaptive Feature Aggregation (AFA) ensembles multiple diffusion models that share the same architecture but different parameters. Specifically, SABW is designed to integrate intermediate features from multiple U-Net denoisers at the block level. Fig. 2 demonstrates an example of ensembling two diffusion models, and each U-Net denoiser contains seven blocks by SABW. The forward process of AFA can be seen in Algorithm 1.
Given $N$ diffusion models to be ensembled, whose U-Net denoisers contain $K$ blocks, let $\boldsymbol{\epsilon}_{\theta_{i}}$ be the $i$-th U-Net denoiser, and $\boldsymbol{\epsilon}_{\theta_{i}}^{(j)}$ be its $j$-th block. For the timestep $t$, it transforms the input features $\mathbf{x}_{t}^{(j)}$ into the output feature by
$\mathbf{y}_{t,i}^{(j)}=\boldsymbol{\epsilon}_{\theta_{i}}^{(j)}(\mathbf{x}_{t}^{(j)},\mathbf{c},t)\ .$ (2)
Note that because of the same architecture of all $\boldsymbol{\epsilon}_{\theta_{i}}$, $\mathbf{y}_{t,i}^{(j)}$ has the same shape, i.e., $\mathbf{y}_{t,i}^{(j)}\in\mathbb{R}^{h_{j}\times w_{j}\times c_{j}}$, where $h_{j}$, $w_{j}$, and $c_{j}$ are the height, width, and channels, respectively. To achieve the output feature for the next block, a block $f_{\varphi}^{(j)}$ of SABW is employed to aggregate $\{\mathbf{y}_{t,i}^{(j)}\}_{i=1}^{N}$,

### v2

@@ Abstract
The success of the text-guided diffusion model has inspired the development and release of numerous powerful diffusion models within the open-source community. These models are typically fine-tuned on various expert datasets, showcasing diverse denoising capabilities. Leveraging multiple high-quality models to produce stronger generation ability is valuable, but has not been extensively studied. Existing methods primarily adopt parameter merging strategies to produce a new static model. However, they overlook the fact that the divergent denoising capabilities of the models may dynamically change across different states, such as when experiencing different prompts, initial noises, denoising steps, and spatial locations. In this paper, we propose a novel ensembling method, Adaptive Feature Aggregation (AFA), which dynamically adjusts the contributions of multiple models at the feature level according to various states (i.e., prompts, initial noises, denoising steps, and spatial locations), thereby keeping the advantages of multiple diffusion models, while suppressing their disadvantages. Specifically, we design a lightweight Spatial-Aware Block-Wise (SABW) feature aggregator that adaptive aggregates the block-wise intermediate features from multiple U-Net denoisers into a unified one. The core idea lies in dynamically producing an individual attention map for each model’s features by comprehensively considering various states. It is worth noting that only SABW is trainable with about 50 million parameters, while other models are frozen. Both the quantitative and qualitative experiments demonstrate the effectiveness of our proposed method. [^1] The code is available at https://github.com/tenvence/afa.
@@ 1 Introduction
Diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020) have progressively become the mainstream models for text-guided image generation (Nichol et al., 2021; Ramesh et al., 2022; Saharia et al., 2022; Rombach et al., 2022; Balaji et al., 2022; Xue et al., 2023; Feng et al., 2023), which treats generation as an iterative denoising task. Recently, the open-source stable diffusion (SD) (Rombach et al., 2022) model has prompted the development and release of numerous powerful diffusion models within the open-source community (e.g., CivitAI [^2] https://civitai.com/). These models are typically fine-tuned on various expert datasets, showcasing diverse denoising capabilities. Leveraging multiple high-quality models to dig out better generations is an important research direction, which has not been extensively studied.
@@
Existing methods leverage multiple diffusion models through the weighted merging of model parameters, which can be called the static method. The weights are usually manually set (e.g., Weighted-Merging [^3] https://github.com/hako-mikan/sd-webui-supermerger and MBW [^4] https://github.com/bbc-mc/sdweb-merge-block-weighted-gui) or automatically searched through enumeration (e.g., autoMBW [^5] https://github.com/Xerxemi/sdweb-auto-MBW). In contrast, model ensembling, which can be called the dynamic method, often uses dynamic strategies to fuse multiple models at the feature level. Unlike ensembling for classification models (Freund & Schapire, 1995) that usually work after the output, ensembling for diffusion models typically needs to work for each block. However, the denoising capabilities of models vary not only at different blocks but also at different spatial locations. As illustrated in Figure 1, we conduct denoising experiments on different prompts with various initial noises. Then, we plot the proportion of wins (i.e., the model with the least error between the predicted noise and the initial noise), for each model in a certain spatial region, to distinguish the denoising capabilities of the models at different states. In other words, different prompts, initial noises, denoising steps, and spatial locations can all have a significant impact on the denoising capabilities of diffusion models. This implies that an adaptive method is needed to ensure that each diffusion model dominates the generation at its strongest states.
@@
In this paper, we propose a novel Adaptive Feature Aggregation (AFA) method, which dynamically adjusts the contributions of multiple models at the feature level by taking into account various states, such as prompts, initial noises, denoising steps, and spatial locations. Specifically, we design a lightweight Spatial-Aware Block-Wise (SABW) feature aggregator that adaptively aggregates the block-level intermediate features from multiple U-Net denoisers into a unified one. The core idea of adaptive aggregation lies in dynamically producing an individual attention map for each model’s features by comprehensively taking into account the various states. A noteworthy aspect of AFA is that only SABW is trainable with about 50 million parameters, while all other models are frozen.
@@ 2 Related Work
Text-Guided Image Synthesis. Early works for text-guided image synthesis leverage Generative Adversarial Networks (GAN) (Goodfellow et al., 2014) conditioned on text (Tao et al., 2022; Xu et al., 2018; Zhang et al., 2021a; Zhu et al., 2019). Based on the success of Transformers (Vaswani et al., 2017), many subsequent works reframe text-guided image synthesis as a sequence-to-sequence task (Zhang et al., 2021b; Ramesh et al., 2021; Ding et al., 2021; Gafni et al., 2022; Yu et al., 2022). Recently, diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020) gradually become mainstream, which treats image generation as an iterative denoising task. By injecting text as a condition into the denoising process, many diffusion-based models achieve significant success in text-guided image synthesis (Nichol et al., 2021; Ramesh et al., 2022; Saharia et al., 2022; Rombach et al., 2022; Balaji et al., 2022; Xue et al., 2023; Feng et al., 2023; Shen et al., 2023a; 2024a; 2024b; Shen & Tang, 2024; Wang et al., 2024; Fu et al., 2024; Shen et al., 2025). Among them, the Latent Diffusion Model (LDM) (Rombach et al., 2022) performs the diffusion and reverse process in the latent space, instead of the pixel space, which largely reduces the computational burden. Among all the implementations of the Latent Diffusion Model, the stable diffusion model (SD) is the most famous one. Along with its public availability, the open-source community achieves tremendous success, with the emergence of many excellent fine-tuned models. These high-quality models are typically fine-tuned on various expert datasets, showcasing diverse denoising capabilities. In this paper, we aim to leverage multiple such models to achieve stronger text-to-image generation.
@@
Model Merging. For vision or language understanding tasks, the effectiveness of parameter merging (Frankle et al., 2020; Wortsman et al., 2022b; Matena & Raffel, 2022; Ilharco et al., 2022b; Li et al., 2022; Don-Yehiya et al., 2022; Jin et al., 2022) can be interpreted from the perspectives of loss landscape (Wortsman et al., 2022a; Ainsworth et al., 2022; Stoica et al., 2023) and task arithmetic (Ilharco et al., 2022a; Ortiz-Jimenez et al., 2023; Zhang et al., 2023). However, merging-based methods have not been widely studied for generation tasks, especially diffusion-based generation. Recently, many intuitive merging-based methods have emerged. One of the popular methods is Weighted-Merging, which manually determines weights to merge each U-Net (Ronneberger et al., 2015) parameter of multiple SD models. Although the simplicity, Weighted Merging coarsely allocates weight to all U-Net blocks. As an improvement, Merge Block Weighted (MBW) allows for manually setting different merging weights for the parameters of distinct U-Net blocks, which provides a more fine-grained merging strategy. Furthermore, to reduce the reliance on manual weights, autoMBW attempts to automate the merging process, which enumeratively selects the optimal combination of MBW weights by an aesthetic classifier. However, autoMBW is constrained by the performance bottleneck of the aesthetic classifier, and the enumerative selection leads to a huge time consumption to find the optimal settings. Such merging methods can be called the static methods. In this paper, we aim to leverage multiple diffusion models by the dynamic methods, i.e., ensembling.
@@
Model Ensembling. Model ensembling is an effective method to achieve better performance (Zhou, 2012), which has been widely applied in various vision understanding tasks, such as classification (Zhao et al., 2005; Rokach, 2010; Yang et al., 2010), regression (Mendes-Moreira et al., 2012; Ren et al., 2015), clustering (Vega-Pons & Ruiz-Shulcloper, 2011). While fewer works focus on the ensembling of generative models, because of the complexity of image space. Vision-Aided GAN (Kumari et al., 2022) focuses on GANs (Goodfellow et al., 2014), which guides the optimization of a target generator by ensembling pretrained models as a loss. MagicFusion (Zhao et al., 2023) focuses on the diffusion models, which fuses the predicted noises of two expert U-Net denoisers to implement specific applications, such as style transferring and object binding. In this paper, we aim to efficiently ensemble multiple diffusion models to achieve general generation improvements [^6] Note that combining the styles/concepts of multiple diffusion models is also a common goal of ensembling/merging, which is not the interest of our work..
@@ 3.1 Preliminary
As a type of generative model, the diffusion model (Sohl-Dickstein et al., 2015; Ho et al., 2020) consists of two processes, which are the diffusion process and the reverse process, respectively. In the diffusion process (i.e., the forward process), the Gaussian noises are iteratively added to degrade the images over $T$ steps until the images become completely random noise. In the reverse process, a trained denoiser is used to iteratively generate images from the sampled Gaussian noise.
@@
$\mathcal{L}_{\text{denoise}}=\mathbb{E}_{\mathbf{x}_{0},\bm{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I}),\mathbf{c},t}\left\|\bm{\epsilon}_{\theta}\left(\mathbf{x}_{t},\mathbf{c},t\right)-\bm{\epsilon}\right\|^{2}\ .$ (1)
Among them, $\mathbf{x}_{t}=\sqrt{\alpha_{t}}\mathbf{x}_{0}+\sqrt{1-\alpha_{t}}\bm{\epsilon}$ is the noisy image at timestep $t\in[1,T]$, where $\alpha_{t}$ is a predefined scalar from the noise scheduler. $\bm{\epsilon}$ is the added noise. $\bm{\epsilon}_{\theta}$ is the denoiser with the learnable parameters $\theta$, which predicts the noise to be removed from the noisy image.
LDM stands out as one of the most popular diffusion models. It performs the diffusion and reverse process in the latent space which is encoded by a Variational Auto-Encoder (VAE) (Kingma & Welling, 2013; Rezende et al., 2014). In LDM, the U-Net structure (Ronneberger et al., 2015) is used for denoiser, which contains three parts of blocks, which are the down-sampling blocks, the middle block, and the up-sampling blocks, respectively. Each down-sampling block is skip-connected with a corresponding up-sampling block. These blocks are composed of Transformer (Vaswani et al., 2017) layers and ResNet (He et al., 2016) layers. In the Transformer layers, the cross-attention mechanism is employed to incorporate conditional textual prompts, which plays a crucial role in text-guided image generation.
Figure 2: Framework of ensembling multiple diffusion models by our AFA method.
@@ 3.2 Adaptive Feature Aggregation
Our proposed Adaptive Feature Aggregation (AFA) ensembles multiple diffusion models that share the same architecture but different parameters. Specifically, SABW is designed to integrate intermediate features from multiple U-Net denoisers at the block level. Figure 2 demonstrates an example of ensembling two diffusion models, and each U-Net denoiser contains seven blocks by SABW. The forward process of AFA can be seen in Appendix A.
Given $N$ diffusion models to be ensembled, whose U-Net denoisers contain $K$ blocks, let $\bm{\epsilon}_{\theta_{i}}$ be the $i$-th U-Net denoiser, and $\bm{\epsilon}_{\theta_{i}}^{(j)}$ be its $j$-th block. For the timestep $t$, it transforms the input features $\mathbf{x}_{t}^{(j)}$ into the output feature by
$\mathbf{y}_{t,i}^{(j)}=\bm{\epsilon}_{\theta_{i}}^{(j)}(\mathbf{x}_{t}^{(j)},\mathbf{c},t)\ .$ (2)
Note that because of the same architecture of all $\bm{\epsilon}_{\theta_{i}}$, $\mathbf{y}_{t,i}^{(j)}$ has the same shape, i.e., $\mathbf{y}_{t,i}^{(j)}\in\mathbb{R}^{h_{j}\times w_{j}\times c_{j}}$, where $h_{j}$, $w_{j}$, and $c_{j}$ are the height, width, and channels, respectively. To achieve the output feature for the next block, a block $f_{\varphi}^{(j)}$ of SABW is employed to aggregate $\{\mathbf{y}_{t,i}^{(j)}\}_{i=1}^{N}$,